# AgenticMath: Enhancing LLM Reasoning via Agentic-based Math Data Generation

### v2

@@ A.2 Training Details
A.3 Evaluation Benchmarks
In-Domain: (i) GSM8K [Cobbe et al., 2021], consisting of grade-school arithmetic word problems that are relatively simple. (ii) MATH [Hendrycks et al., 2021], a large-scale dataset of competition-level problems that are significantly more challenging.
Out-of-Domain (OOD): (i) CollegeMath [Tang et al., 2024], with 2,818 college-level problems drawn from nine textbooks across seven domains (e.g., linear algebra, differential equations), designed to test generalization to complex mathematics. (ii) DeepMind-Mathematics [Saxton et al., 2019], a collection of 1,000 problems covering a national school curriculum (up to age 16), assessing basic reasoning across varied types. (iii) OlympiadBench-Math [He et al., 2024], providing 675 Olympiad-level problems (English text-only subset) targeting the most challenging reasoning tasks. (iv) TheoremQA [Chen et al., 2023], consisting of 800 problems that require applying mathematical theorems across mathematics, physics, and engineering, testing theoretical reasoning in STEM.
A.4 Examples of Problem Revision
A.4.1 Examples of Problem Revisions from GSM8K
@@
A.4.2 Examples of Problem Revisions from MATH
@@ A.5 Agent prompt
Example 3: Problem Review Prompt As a mathematics quality checker, your task is to rigorously assess whether a given mathematical question is high-quality and provide rewrite suggestions: 1. Clarity & Grammar (1–5): The question must be grammatically correct, precisely phrased, and easy to understand. It should avoid ambiguity in wording or phrasing. 2. Logical Coherence & Completeness (1–5): All elements of the problem (e.g., given information, constraints, relationships, objectives) must be logically interconnected and sufficient. The problem should present a clear, sequential path for reasoning, without missing information required for the specified solution approach. 3. Mathematical Validity & Solvability (1–5): The problem must be fundamentally a mathematics problem, with all its premises and conditions being *mutually consistent* and *mathematically sound*. It must lead to a *unique, solvable numerical or analytical answer* that adheres to all mathematical rules and specified ranges (e.g., probabilities summing to 1, valid geometric properties, real number solutions). If any condition leads to a mathematical contradiction or an impossible/undefined solution (e.g., total probability >> 1 after adjustments, an equation with no valid solution within given constraints), this criterion rates very low, and the exact mathematical inconsistency must be pinpointed. Avoid open-ended or non-mathematical questions. ** Scoring Guidelines **: - Please rate the sample on a scale from 1 to 5 for each criterion, and return an overall rating on a scale from 1 to 5, where a higher score indicates higher level of quality. Rephrased question: {rephrased_question} **Output Requirements** Respond in the following plain-text format **only** (do not include JSON or any additional commentary): ###thought### <Analytical reasoning addressing each criterion sequentially, especially for rephrased_question > ###rating_score### ["<Clarity & Grammar score >>", "<Logical Consistency score >>", "<Mathematical Relevance & Solvability score >>"] ###suggestions### ###Specific improvement 1### <Specific improvement 1 > ###Specific improvement 2### <Specific improvement 2 > …more improvements if needed… Noice: - "rating_score" represents evaluate score of Rephrased question. - when generate "suggestions", please give more details and reasons for each improvement.

### v3

@@ A.11 Related work
Figure 3: t-SNE Semantic Distribution.
A.12 Examples of Problem Revisions from GSM8K
@@
A.13 Examples of Problem Revisions from MATH
@@ A.14 Agent prompt
Example 3: Problem Review Prompt As a mathematics quality checker, your task is to rigorously assess whether a given mathematical question is high-quality and provide rewrite suggestions: 1. Clarity & Grammar (1–5): The question must be grammatically correct, precisely phrased, and easy to understand. It should avoid ambiguity in wording or phrasing. 2. Logical Coherence & Completeness (1–5): All elements of the problem (e.g., given information, constraints, relationships, objectives) must be logically interconnected and sufficient. The problem should present a clear, sequential path for reasoning, without missing information required for the specified solution approach. 3. Mathematical Validity & Solvability (1–5): The problem must be fundamentally a mathematics problem, with all its premises and conditions being *mutually consistent* and *mathematically sound*. It must lead to a *unique, solvable numerical or analytical answer* that adheres to all mathematical rules and specified ranges (e.g., probabilities summing to 1, valid geometric properties, real number solutions). If any condition leads to a mathematical contradiction or an impossible/undefined solution (e.g., total probability >> 1 after adjustments, an equation with no valid solution within given constraints), this criterion rates very low, and the exact mathematical inconsistency must be pinpointed. Avoid open-ended or non-mathematical questions. ** Scoring Guidelines **: - Please rate the sample on a scale from 1 to 5 for each criterion, and return an overall rating on a scale from 1 to 5, where a higher score indicates higher level of quality. Rephrased question: {rephrased_question} **Output Requirements** Respond in the following plain-text format **only** (do not include JSON or any additional commentary): ###thought### <Analytical reasoning addressing each criterion sequentially, especially for rephrased_question > ###rating_score### <Clarity & Grammar score >, <Logical Consistency score >, <Mathematical Relevance & Solvability score > ###suggestions### ###Specific improvement 1### <Specific improvement 1 > ###Specific improvement 2### <Specific improvement 2 > …more improvements if needed… Noice: - "rating_score" represents evaluate score of Rephrased question. - when generate "suggestions", please give more details and reasons for each improvement.